# Blockchain Address Poisoning

### v1

@@ Appendix F UI Design of wallets and chain scanners
To discover how victims see the attacks on their wallets and chain scanners, we first implement address poisoning transfers and attack our own accounts (i.e., attacking our accounts) on the Ethereum mainnet. While our attack uses minimal blockchain resources (three poisoning transfers), conducting our attack on the mainnet is critical because of the significant UI difference from the testnet. Specifically, we conduct our experiment on Dec. 20th, 2024, and used 159,364 gas for three poisoning transfers which consumes about 1% of one block (0.00014% of gas usage for a day).
We generate a lookalike address $L$ that matches the first six and the last 5 characters to $R$. We follow to $AC$ and $CT$ implementations from the attackers who have uploaded their contracts’ source codes on Etherescan and perform three poisoning transfers on our victim account $V$. Our performed attack can be found in $V$ 0xB6e84DF1cE401117C450221ccc6EF502cb0e2284.
@@
First, we check the victim’s UI on blockchain wallets (as of Jan. 2nd, 2025). Based on the wallet market share [25], we study MetaMask, Trust Wallet, and Phantom. Regarding the address abbreviations (first, last), Metamask is (4,4), Trust Wallet is (2, 4), and Phantom is (4, 4). Trust Wallet and Phantom only display tiny transfers but not zero-value or counterfeit token transfers. Metamask does not appear to show any ERC-20 token transfer history in the first place. When initiating the transfers, users no longer have options to select the recipient from “recently used addresses” but manually specify it on their own. The fact suggests that victims may rely more on blockchain scanning services to find the recipients.
@@
Second, we check the major blockchain scanning services: Etherscan, Ethplorer, OkLink, and Arkm (as of Jan. 2nd, 2024). Table 8 compares the UI of those blockchain scanners (tiny: T, zero-value: Z, and counterfeit token: C). The level of address abbreviations significantly differs by each service. Etherscan equally shortens the beginning and the end while others significantly abbreviate one side (or no abbreviation at all). Since the attackers mostly focus on the first and last characters now, displaying one side might be effective. Next, we check the poisoning transfers. By default, most services (but Ethplorer) successfully hide zero-value transfers while they fail to detect tiny transfers. By changing the setting, Etherscan and OkLink additionally show zero-value, and Arkm displays counterfeit token transfers. It is generally easier for them to hide tiny than zero-value (i.e., just detecting the value of 0) while they must implement some heuristics to detect tiny or counterfeit token transfers [^22] Etherscan successfully detects all poisoning transfers as of our visit on Jan. 13th, 2025.. Note that hiding zero-value transfers hinders the usability if users use zero-value transfers for (benign) testing purposes. Most services show token icons or give a warning to $CT$ (based on their list of $AC$), which would be helpful for users to distinguish them.
Table 8: UI comparison of blockchain scanners
Addr. abbr. Poisoning transfers Token icon Default Setting change Etherscan 8, 9 T C Z Yes Ethplorer No abbr. T, Z, C No OKLink 4, 11 T, C Z Yes Arkm 36, 0 T C Yes
Appendix G Payoff transfers
Table 9 presents the 30 largest payoff transfers on the Ethereum blockchain and BSC as examples for readers to investigate. These transactions are the final step of the attack, in which victims mistakenly send funds to the attackers. We verify the authenticity of these attacks discovered by our algorithm on Etherscan (BSCscan) through the following steps:
  1. 1.
    In the transaction page, get the sender (i.e., from) of the Transfer event
  2. 2.
    Go to the sender (victim)’s page and examine the Token Transfer (ERC-20) tab.
  3. 3.
    Follow the prompt and change the site settings to disable “Ignore Tokens with Poor Reputation” and “Zero-Value Token Transfers” options.
  4. 4.
    Find the payoff transfer transaction and verify that there are attack attempts in earlier blocks. For example, in the first case of Ethereum, a transfer of 10,000,000 counterfeit tokens from the victim to the (attacker’s) lookalike address took place in block 17,778,047.
  5. 5.
    Find the authentic token transfer events from the victim to the intended address that share similar characters to the lookalike address. For example, in the first case of Ethereum, a transfer of 10,000,000 USDT (Tether) is sent from the victim to the address 0xa7B4BAC8f0f9692e56750aEFB5f-
    6cB5516E90570
Table 9: Largest payoff transfers on Ethereum and Binance Smart Chain. USD Phished is the monetary loss in this single transaction, Recipient is the attacker’s lookalike address, Intended recipient is the address that the victim interacted with, Matched is the number of characters matched between the lookalike and the intended recipient, $a+b$ indicates that the lookalike address matched the first $a$ characters and the last $b$ characters of the intended address.
(a) Ethereum blockchain
USD Phished Transaction Hash Recipient (attacker) Intended recipient Matched
20,000,000 0x08255ca0e42a872559437141fa46980e66d907f7668922467d67515b1ebb4b7f 0xa7Bf48…E90570 0xa7B4BA…E90570 3 + 6
3,800,000 0x2dea87ff34b0d4712d59e582e810d1a8396e469ce7830a61b9521759b28221f1 0x61b697…074c5C 0x61b9d3…174c5c 3 + 5
3,554,610 0x11ac26acce0620c50731a94b528419771211779903d6917b05f2bd90e4a6476c 0x1cbB23…9B758a 0x1CbB23…9B758a 7 + 6
2,030,000 0x0911bd6713493e9ab75ef82cc909114218996f0e717b81b71d7bf29fa06e1622 0x74C9bd…60E1cA 0x74C32c…50E1cA 3 + 5
2,000,000 0x48e591d562a5098527c0de850ba44ce21014726434184f693707f62877a16994 0xbb2EDb…619455 0xBB217c…219455 3 + 5
2,000,000 0x5ae3147442cb9cba98b8b3dc1548a7c27dab29847ca2c126059ca806ad4479c9 0xC7B14b…B33A8f 0xc7b57d…533a8F 3 + 5
1,999,000 0x7cae29b7d215ef67fa183d4455760fc6e8698bfb2c5f1744dec4749d0c4b88c7 0xC7B14b…B33A8f 0xc7b57d…533a8F 3 + 5
1,660,000 0x86d39a9fd0e4078223179e6f59d2651963edb07899efc9558c43a0a338fb40a5 0xEA0d13…CBeb14 0xEa0ddb…FcEb14 4 + 4
1,499,990 0xf9948a832bfc217451f056fb3dc2df258c6ad525d7c3c184e07eeeee3df7ed09 0x9339e7…1F4126 0x9339CF…3A4126 4 + 4
1,200,000 0x371f26eb4b08b37d69b280fa5ea06208750c4bc9cf3258950d63c842f9f01154 0xCbA796…34c994 0xCbA796…34c994 8 + 6
1,045,150 0xa7be1b408b17853c3f9de8cef05371616d154b646e0f7c556feb72dc87ca9676 0x73435A…ca79F7 0x734659…Ca79F7 3 + 6
1,000,000 0xb97df9feb55c3849da4551f43995ec0f0fd78ff0e3d28f337934dbaba154bc11 0x80D707…4dbeA2 0x80D713…f4beA2 4 + 4
960,000 0x5b03c7356fe51575bd4fd190cb83841ec7398d8531dd34d21d89fdc0f776afd5 0x946C8e…f98Cd6 0x9462B5…f98CD6 3 + 6
865,884 0x71eb719e26a31dbb9b05d38b8a36d73537af35e3a5e0d2ccf4bbb6accbd8cd36 0x55B714…Ba9d79 0x55be3c…Ca9d79 3 + 5
849,996 0xb09f4c9fe09e6bbeb3abfdbed21a0403a2d8d2b6313ed1cbcdbd620a1f748fb2 0xEB4034…84683d 0xEb4034…84683D 8 + 7
800,158 0xaeb6f1b730e65ddb05c84d71af8c370dead57c4d1aecafc7e2369fde50875ae6 0x0Cee6B…e70EbF 0x0CEE33…3f0EbF 4 + 4
749,900 0xecd48ffd8d6a09f1ca18e7acbaad4ff4f70116d70fa65dab842a7cd1651df8a0 0xd7F373…7B7403 0xd7F38D…7B7403 4 + 6
719,000 0x1576beff54a53f0ef0d881a6242edf4539d0964b5b2431f9e1f08fe781b42de3 0x9887E5…D40ea3 0x9887F3…740ea3 4 + 5
710,000 0x66096743f07f2f0d49818f5e4de28b98bbf3ac6bfcf665fa48225f97260be6b1 0x949D0D…ef861B 0x949954…58861B 3 + 4
699,990 0xb2d6e74cba922beeae6dfd9f8a832372ae1d4d5314c0d1505dda13911e840316 0xC7BA77…26b438 0xc7BD59…F5B438 3 + 4
618,000 0x273166dfae3d85e8297a0c1353d8d9bc72ab7b021a61797cd30b8d1871f59fea 0x55Eb2A…A83159 0x55e5Cf…083159 3 + 5
600,000 0xf997e3b1fcba63c52815dbc63d4364691a79ef65bb578957f182f54753ef59ca 0xE07222…6adFeA 0xE07C8D…6AdFea 3 + 6
561,035 0x65f005042adff7a66fd5e317459987367dcfbae3246a50dbde88d1f364c4d9a5 0x368595…c3aFd3 0x3683a1…C3aFD3 3 + 7
541,900 0x82be73572215148ba7c03e984b1361188b098b7ceedf1aaac2a2574ec3f52fa0 0x35D6CD…4B0305 0x35D3c7…4b0305 3 + 6
541,000 0x7b5b69fad190d0e2b233d1b5bc0fb4a92178b72a737b3f2041a438a8bcd8ecad 0x60855B…30D466 0x6085fd…c6D466 4 + 4
500,962 0xea71be058fcd162620d95d04b2d5acbec953cedbb5ccb5339e6d43198b4b026e 0x539592…0DDf0b 0x539683…a6Df0B 3 + 4
500,000 0x8e452c91d99185a0ba325b2ea3dbb276d09f4b2a6dc6772d615acd48237a3148 0x324AFa…80C28F 0x32494C…C0c28F 3 + 5
500,000 0xd4bc53f71576fc1b78dd56145e42a127d4109567e7d04c3033e59d9ce9da502c 0xB08cfc…A74a2F 0xb08D97…574A2f 3 + 5
500,000 0xe68134277191eae950ca698b6651ac7c1942c6c4578eb53a98379c7deceb3f33 0x140588…b0Cb01 0x140b09…25cB01 3 + 4
500,000 0x78cfaa6c26a06444561397b7b02a426b1e3250b1eaf7e051e91b9137a42fdf86 0xC6BCCb…Bf0115 0xc6B7A8…BF0115 3 + 6
(b) Binance Smart Chain
USD Phished Transaction Hash Recipient (attacker) Intended recipient Matched
279,489 0xa0880a04b9f4dd114eb6e5b457a4b6f260c1e1444eef350b62cdc5f1cba8b0ca 0x7e4d7a…a7f223 0x7e4D54…dcf223 4 + 4
150,000 0xba144f4c4542434bc1ebc2d70030a89f46e81c993842bf79b43ad453cd446d3e 0x112C71…C4B08a 0x112784…24B08a 3 + 5
144,876 0xc11c5930465a2c33a479b7df63eaa15610dd9ca78c70f788d0f69cfe5e96960d 0x25f566…b02c50 0x25f5e8…b02c50 4 + 6
133,000 0x266647bc5190cd516f791399ac61b10e876ad75dcc0c6fabb77ccc80f7ead038 0xe72a80…49dc98 0xe72dee…f7dc98 3 + 4
100,000 0x62798e183339f5ed70239a84b4463b7e48f497e1ef0e23ee82b78087195462d4 0xD643f1…eAF6fB 0xd64329…24F6fB 4 + 4
97,250 0x9a4081519e99e2e6fe6add185da7333d44eb28cd1c70df36a32ff8648940cbf5 0x8AC739…5cEB7B 0x8ac7CE…76eb7b 4 + 4
91,680 0x12505feedc126def4e118a2b8e970766795d45201a79dbe45f4b2b9a41391375 0x5ebefa…8De436 0x5EBe2a…EDe436 4 + 5
86,180 0xa070b46843b57ff4e529cbe7aabb67c99cb36da58510106bb464d1f039e7b3db 0x2556be…757cF1 0x255EC6…A57cF1 3 + 5
72,671 0x4e7dd743e32d51f6f8d0705d20542a15a9a8b40a80e707ad1e1f18147033b81b 0x6c9351…D01114 0x6c93Ce…EC1114 4 + 4
70,006 0x13aecc1f514cc95db38a31ada1c9827124aa87360dc11cb1b3dfd7fe0d2499e0 0x2f9233…2BC04D 0x2f9276…cbc04d 4 + 5
60,000 0x521bb785beaac9b08cc35d57096a40d28dc7e3e697fe22572a0431681770cc04 0x1616bB…4d1800 0x161c2C…6d1800 3 + 5
50,111 0xe5576fff71817e990959b6947aa54866888953e2912d3047620b27d0ac766350 0xf06e0a…b8b610 0xf06e59…29b610 4 + 4
50,094 0x09c21db09c5e92e68bf01eb46374796a9a3c16a0b42df9ef2217020f2f8e2352 0x5c3415…82E06f 0x5c3464…a2E06F 4 + 5
50,000 0x6108401848776037926010c09b848f00ec5e04e3c4075b72437575544e4c0cd9 0x809799…4A2d4c 0x8097dc…ba2D4C 4 + 5
50,000 0xb4a277c05232537bbb1dd2aa1f04c8dced24fbd5be125c65fe0498db3f4b837d 0x1487c9…f08B32 0x148799…b08b32 4 + 5
50,000 0xedfe46269453c28ff4d70f736ca73664ca032bbc65a67462b3d35ea178e47171 0xC4d43e…71d48D 0xc4D89C…81D48D 3 + 5
46,000 0xe4335fecccba5eec21dbaa18d71cdb4d1c56056ee38c8a30b2a57aaccf4c4faa 0xcc3af0…97298f 0xcc3afe…fe298f 5 + 4
40,008 0x230ef03a64905d8e0c9a074fde33c0fb993ad536b3852129c36efe485409ad69 0x458728…dF9961 0x4587C1…629961 4 + 4
39,995 0x5ac785bf1bc5273ca3183471bafce3e1a301afc969f3e5eb26e83f249cca70b4 0x73A31C…E5A703 0x73af35…35a703 3 + 5
37,000 0x41c732fa968e3816c69ffaebc849233b84a863695344f24328891654b1230edf 0x52C990…0c1961 0x52c908…fC1961 4 + 5
35,000 0xd96aa7bfbd32720c88ecfddaec21af640a65cb858ddb735d23b5d6b1fae2b1bc 0xe6287d…08d0EA 0xe62d77…f8D0ea 3 + 5
32,000 0x7cee39ef7f56a08b88346938540be29d58249a33b1b53a1166203d7c467e21c4 0xF632D3…294ee0 0xf63078…194EE0 3 + 5
30,060 0x53eb5af3962cf511e51b6369071494118460cc11e24234a3551478f3415e0a47 0x4245B2…782880 0x424276…782880 3 + 6
30,000 0x6bc9148d278302517d49f7602f5c8fd0ef516ea8563996910db846ab029e6994 0x8b8ea3…7A8F9F 0x8B808A…5a8f9f 3 + 5
30,000 0xa9c015c3b41dd449266cb67fb8c514aadc501fac88a1c49e06d6b51b0c0a9953 0x532D78…C1540f 0x532Dc7…71540f 4 + 5
30,000 0x6462aaf80cc01700133399d6086cc0381f72ae48b52cdd85529fd7ec45d19fc9 0x2FD1D3…5cc8B8 0x2fd1Da…cCC8B8 5 + 5
30,000 0x6ef40292302c45c1a0a1756b9b431e4b08f02118ca78c12892377c52e5a84796 0xBB3e51…c3aEAF 0xBb3fd3…93aEAF 3 + 5
29,377 0x74123f8aef47435e94a4c2291df65bdcf1d6735de97e05fb84974c48a207983d 0xE15D59…5bA949 0xE15f47…6ba949 3 + 5
28,650 0x1e56b900524521ef21e82392ff8992675a58f0af74a8ddf476c96876533cd5c2 0xaF64F8…Cf01d0 0xaF6859…3f01d0 3 + 5
25,666 0x60b3ca5895303d08fe83288419e805de7cd078b1778ba7ee30ceafdbac170601 0x4A2274…7b57A0 0x4a22aE…bB57a0 4 + 5

### v3

@@ Appendix G UI Design of wallets and chain scanners
To discover how victims see the attacks on their wallets and chain scanners, we first implement address poisoning transfers and attack our own accounts (i.e., attacking our accounts) on the Ethereum mainnet. While our attack uses minimal blockchain resources (three poisoning transfers), conducting our attack on the mainnet is critical because of the significant UI difference from the testnet. Specifically, we conduct our experiment on Dec. 20th, 2024, and use 159,364 gas for three poisoning transfers, which consumes about 1% of one block (0.00014% of gas usage for a day). We generate a lookalike address $L$ that matches the first six and the last four characters to $R$. We follow to $AC$ and $CT$ implementations from the attackers who have uploaded their contracts’ source codes on Etherscan and perform three poisoning transfers on our victim account $V$. Our attack can be found in $V=$0xB6e84DF1cE401117C450221ccc6EF502cb0e2284.
@@
First, we check the victim’s UI on blockchain wallets (as of Jan. 2nd, 2025). Based on their wallet market share [26], we study MetaMask, Trust Wallet, and Phantom. All three services abbreviate addresses by only showing the first $a$ and last $b$ characters. Metamask uses $(a,b)=(4,4)$, Trust Wallet uses (2, 4), and Phantom uses (4, 4). Trust Wallet and Phantom only display tiny transfers but hide zero-value or counterfeit token transfers. MetaMask does not seem to show any ERC-20 token transfer history in the first place. When initiating transfers, users no longer have options to select the recipient from “recently used addresses” but manually specify it on their own. The fact suggests that victims may rely more on blockchain scanning services to find the recipients.
@@
Second, we check four major blockchain scanning services: Etherscan, Ethplorer, OkLink, and Arkham (as of Jan. 2nd, 2025). The level of address abbreviation significantly differs between services. Etherscan elides the middle part of the address while keeping the same number of characters at the beginning and at the end, while others significantly abbreviate one side (or do not abbreviate at all). Since attackers mostly focus on collision on the first and last characters, displaying one side might be effective. Next, we check the poisoning transfers. Most services (but Ethplorer) successfully hide zero-value transfers (by default) but fail to detect tiny transfers. By changing settings, Etherscan and OkLink can show zero-value transfers, and Arkham displays counterfeit token transfers. It is generally easier for a service to flag zero-value transfers than tiny/counterfeit token transfers (by just observing their value); scanning services must implement more advanced heuristics to detect tiny or counterfeit token transfers [^28] Etherscan successfully detects all poisoning transfers as of our visit on Jan. 13th, 2025.. Hiding all zero-value transfers also hinders usability if users use zero-value transfers for (benign) testing purposes. Most services show token icons or give a warning to $CT$ (based on their list of $AC$), which should be helpful to users.